Mathematics 242 – Principles of Analysis Definitions Quiz 8 – April 11, 2014

Your Name:

A) (5) Let f be a bounded function on the interval [a, b]. What does it mean for f to be integrable on [a, b]?

B) (5) True/False and explain: If f has domain equal to the interval (a, b) and is differentiable with f′(x) > 0 for all x ∈ (a, b), then f is a one-to-one function.